OpenStudy (anonymous):

Check my work? "♥" marks what i chose. If y varies inversely with x and y = 6 when x = 3, then what is y when x = 2? 9 12 18 ♥ 36

jimthompson5910 (jim_thompson5910):

"y varies inversely with x" means that y = k/x for some constant k plug in y = 6, x = 3 y = k/x 6 = k/3 6*3 = k 18 = k k = 18 So the equation is y = 18/x Now plug in x = 2 y = 18/x y = 18/2 y = 9

OpenStudy (anonymous):

Ohh so i was somewhat wrong. thank you :]

jimthompson5910 (jim_thompson5910):

you're welcome